# Constrained inhomogeneous spherical equations: average-case hardness

### v1

@@ Abstract.
In this paper we analyze computational properties of the Diophantine problem (and its search variant) for spherical equations $\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=1$ (and its variants) over the class of finite metabelian groups $G_{p,n}={\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$, where $n\in{\mathbb{N}}$ and $p$ is prime. We prove that the problem of finding solutions for certain constrained spherical equations is computationally hard on average (assuming that some lattice approximation problem is hard in the worst case).
Keywords. Spherical equations, finite groups, semidirect products, metabelian groups, average case complexity, hash function family, group-based cryptography.
@@ 1.1. Equations in groups
Let $F=F(Z)$ denote the free group on countably many generators $Z=\{z_{i}\}_{i=1}^{\infty}$. For a group $G$, an equation over $G$ with variables in $Z$ is an equality of the form $W=1$, where $W\in F\ast G$. If $W=z_{i_{1}}g_{1}\cdots z_{i_{k}}g_{k}$, with $z_{i_{j}}\in Z$ and $g_{j}\in G$, then we refer to $\{z_{i_{1}},\ldots,z_{i_{k}}\}$ as the set of variables and to $\{g_{1},\ldots,g_{k}\}$ as the set of constants (or coefficients) of $W$. We occasionally write $W(z_{1},\ldots,z_{k})$ or $W(z_{1},\ldots,z_{k};g_{1},\ldots,g_{k})$ to indicate that the variables in $W$ are precisely $z_{1},\ldots,z_{k}$ and (in the latter case) the constants are precisely $g_{1},\ldots,g_{k}$. A solution for an equation $W(z_{1},\ldots,z_{k})=1$ over $G$ is an assignment for variables $z_{1},\ldots,z_{k}$ that makes $W=1$ true.
@@ 1.2. Constrained equations
Constrained equations in groups have received very little attention; only certain equations in free groups have been studied. For instance, in [9] V. Diekert demonstrated decidability and ${\mathbf{PSPACE}}$-hardness of the following problem: given an arbitrary equation
$w(x_{1},\dots,x_{n},a_{1},\dots,a_{m})=1$
in a free group $F_{m}$ and a list $H_{1},\dots,H_{n}$ of regular subsets of $F_{m}$, decide whether there exists a solution of this equation satisfying $x_{i}\in H_{i}$. V. Durnev showed in [11] that for a free group $F=F(a,b)$ there exists an equation of type
$w(x^{k},x_{1},\dots,x_{n})=[a,b]$
for which there is no algorithm to recognize if, for a given $k$, it has a solution satisfying $x_{1},\dots,x_{n}\in[F_{2},F_{2}]$.
@@ 1.3. Spherical equations
(1) $\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}$ $=1$ $m\geq 1,$
(2) $\prod_{i=1}^{g}[x_{i},y_{i}]\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}$ $=1$ $g\geq 1,m\geq 0,$
(3) $\prod_{i=1}^{g}x_{i}^{2}\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}$ $=1$ $g\geq 1,m\geq 0.$
@@ 1.5. Model of computation
2. Preliminaries: knapsack problems in groups
@@
(5) $g=g_{1}^{\varepsilon_{1}}\cdots g_{k}^{\varepsilon_{k}}$
@@
By [31, Proposition 2.5], computational properties of ${\mathbf{SSP}}$ do not depend on the choice of a finite generating set $X$ and, hence, the problem can be abbreviated as ${\mathbf{SSP}}(G)$. The same paper, [31], provides a variety of examples of groups with ${\mathbf{NP}}$-complete (or polynomial time) subset sum problems.
Consider the infinitely generated group ${\mathbb{Z}}_{3}^{\omega}$ whose elements can be formally viewed as functions $f\colon{\mathbb{N}}\to{\mathbb{Z}}_{3}$ with finite support. For algorithmic purposes, we assume that elements of ${\mathbb{Z}}_{3}^{\omega}$ are encoded by finite ternary strings (as in [35, Section 4]). [36, Proposition 2.1] proves that ${\mathbf{SSP}}({\mathbb{Z}}_{3}^{\omega})$ is ${\mathbf{NP}}$-complete, which can be reformulated as follows.
Proposition 2.1 ([36, Proposition 2.1]).
@@ 3.2. Short integer solution problem
(7) $\left\{\begin{array}[]{l}A{\overline{x}}\equiv_{p}{\overline{0}},\\ {\overline{x}}\mbox{ is short and nontrivial},\\ \end{array}\right.$
@@
Randomized ${\mathbf{SIS}}$ problem. For a matrix $A\leftarrow{\mathbb{Z}}_{p}^{n\times m}$ sampled uniformly randomly, find ${\overline{x}}\in{\mathbb{Z}}_{p}^{m}$ such that $A{\overline{x}}={\overline{0}}$, and ${\overline{x}}$ is short and nontrivial.
Theorem 3.1 (Goldreich, Goldwasser, Halevi, [16, Theorem 1], cf. Ajtai, [3, Theorem 1]).
@@
(8) $n\log(p)<m<\tfrac{p}{2n^{4}}\ \mbox{ and }\ p=O(n^{c})\mbox{ for some }c>0,$
@@
Randomized ${\mathbf{ISIS}}$ problem. For a matrix $A\leftarrow{\mathbb{Z}}_{p}^{n\times m}$ and a vector ${\overline{y}}\in{\mathbb{Z}}_{p}^{n}$ sampled uniformly randomly, find ${\overline{x}}\in{\mathbb{Z}}_{p}^{m}$ such that $A{\overline{x}}={\overline{y}}$, and ${\overline{x}}$ is short.
Corollary 3.2.
The statement of Theorem 3.1 holds for the randomized ${\mathbf{ISIS}}_{\{0,1\}}$ problem.
Proof.
@@
This gives a PPT algorithm that solves ${\mathbf{SIS}}_{\{0,1\}}$ with probability at least $n^{-c_{0}}$. ∎
4. Spherical equations over ${\mathbb{Z}}_{p,n}\rtimes{\mathbb{Z}}_{p}^{\ast}$
@@
Proof.
Straightforward verification. ∎
4.1. Spherical equations over ${\mathbb{Z}}_{p,n}\rtimes{\mathbb{Z}}_{p}^{\ast}$: generic-case
Here, we investigate generic-case hardness for spherical equations, i.e., hardness of a “typical” equation, see [32, Chapter 10] for basic definitions of generic-case complexity.
Proposition 4.2.
@@ 4.1. Spherical equations over ${\mathbb{Z}}_{p,n}\rtimes{\mathbb{Z}}_{p}^{\ast}$: generic-case
is a solution that can be computed in polynomial time. ∎
@@
that converges to 00 exponentially fast. The obtained bound $\frac{1}{2^{s/2}}$ does not depend on $p$ or $n$. Therefore, the bound works on the whole ${\mathcal{I}}_{s}$. ∎
Lemma 4.4.
@@
Proof.
By Lemma 4.3, we may assume that $m\geq s/2$, in which case
@@
which converges to $1$ exponentially fast as $s\to\infty$. ∎
Corollary 4.5.
@@
Proof.
Follows from Proposition 4.2 and Lemma 4.4
@@
(9) $C_{p,n}=\left\{\,({\overline{c}},1)\;\middle|\;{\overline{c}}\in{\mathbb{Z}}_{p}^{n}\,\right\}.$
From now on we assume that all $c_{i}\in C_{p,n}$. In that case condition (S1) of Lemma 4.1 is trivially satisfied and condition (S2) translates into the following:
(10) $\exists\alpha_{i}\in{\mathbb{Z}}_{p}^{\ast}\ \ \mbox{ s.t. }\ \sum_{i=1}^{m}\alpha_{i}^{-1}{\overline{y}}_{i}\equiv_{p}{\overline{0}}.$
@@
4.2. Spherical equations over ${\mathbb{Z}}_{p,n}\rtimes{\mathbb{Z}}_{p}^{\ast}$: worst case
@@
(11) $\prod_{i}^{m}z_{i}^{-1}c_{{\overline{v}}_{i}}z_{i}=c_{{\overline{v}}^{\prime}}.$
Proposition 4.6.
@@
Proof.
By Proposition 4.6, $I\to E_{I}$ is a many-one polynomial-time reduction of ${\mathbf{SSP}}({\mathbb{Z}}_{3}^{n})$ to spherical equations over ${\mathbf{SSP}}({\mathbb{Z}}_{3,n}\rtimes{\mathbb{Z}}_{3}^{\ast})$. By Proposition 2.1, ${\mathbf{SSP}}$ is ${\mathbf{NP}}$-complete for groups $\{{\mathbb{Z}}_{3}^{n}\}_{n=1}^{\infty}$. Hence the result. ∎
@@ 5. Constrained spherical equations: average-case hardness
In this section we discuss spherical equations over the groups $G_{p,n}$ with constraints on values of $z_{i}$’s. In full generality the problem can be formulated as follows. Given a group $G\in{\mathcal{G}}$, a spherical equation $\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=1$ over $G$, and subsets $Z_{1},\dots,Z_{m}\subseteq G$, find $z_{1},\dots,z_{m}\in G$ satisfying
@@
It is not hard to prove that ${\mathbf{CISE}}$ for groups $G_{p,n}$ is hard in the worst case. In fact, for $p=3$, the proof of Proposition 4.6 works as is, with constraints $Z_{i}=\{({\overline{0}},1),({\overline{0}},2)\}$ (so chosen sets $Z_{i}$ effectively define no constraints). Using these constraints we can easily achieve the same result for any odd value of $p$.
Theorem 5.1.
@@
Proof.
For a given instance ${\overline{v}}_{1},\dots,{\overline{v}}_{m},{\overline{v}}\in{\mathbb{Z}}_{p}^{n}$, call it $I$, of ${\mathbf{SSP}}({\mathbb{Z}}_{p}^{n})$, as in the proof of Proposition 4.6, construct a constrained inhomogeneous spherical equation $E_{I}$ over ${\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$
@@
and observe that $I$ is a positive instance of ${\mathbf{SSP}}({\mathbb{Z}}_{p}^{n})$ if and only if $E_{I}$ has a solution. This gives a polynomial time reduction from ${\mathbf{SSP}}$ over groups $\{{\mathbb{Z}}_{p}^{n}\}_{n=1}^{\infty}$ to decidability of constrained inhomogeneous spherical equations over ${\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$. ∎
@@
(12) $\left\{\begin{array}[]{l}\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=c\ \ \mbox{ with }c_{i},c\in C_{n}\\ z_{i}=({\overline{0}},1)\mbox{ or }({\overline{0}},2^{-1}).\end{array}\right.$
@@
(13) $\left\{\begin{array}[]{l}\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=c\ \ \mbox{ with }c_{i},c\in C_{n}\\ z_{i}=({\overline{0}},1),({\overline{0}},2^{-1}),\mbox{ or }({\overline{0}},3^{-1}).\end{array}\right.$
@@
This establishes a one-to-one correspondence between ${\mathbf{ISIS}}_{\{0,1\}}$ and ${\mathbf{CISE}}_{\{1,2\}}$ that translates the uniform distribution for the instances of ${\mathbf{ISIS}}_{\{0,1\}}$ to the uniform distribution for the instances of ${\mathbf{CISE}}_{\{1,2\}}$ (the element ${\overline{v}}^{\prime}\in C_{n}$ is uniformly distributed because ${\overline{y}}$ uniformly distributed). A similar reduction can be used to establish a one-to-one correspondence between ${\mathbf{SIS}}_{\{-1,0,1\}}$ and ${\mathbf{CISE}}_{\{1,2,3\}}$. ∎
Remark 1.
The one-to-one correspondence established in the proof of Theorem 5.2 also implies that uniformly chosen instances of ${\mathbf{CISE}}_{\{1,2\}}$ and ${\mathbf{CISE}}_{\{1,2,3\}}$ have solutions with probability approaching $1$ as $n\to\infty$.
Remark 2.

### v2

@@ Abstract.
In this paper we analyze computational properties of the Diophantine problem (and its search variant) for spherical equations $\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=1$ (and their variants) over the class of finite metabelian groups $G_{p,n}={\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$, where $n\in{\mathbb{N}}$ and $p$ is prime. We prove that the problem of finding solutions for certain constrained spherical equations is computationally hard on average (assuming that some lattice approximation problem is hard in the worst case).
Key words and phrases:
Spherical equations, finite groups, semidirect products, metabelian groups, average case complexity, hash function family, group-based cryptography.
@@ 1.1. Equations in groups
Let $F=F(Z)$ denote the free group on countably many generators $Z=\{z_{i}\}_{i=1}^{\infty}$. For a group $G$, an equation over $G$ with variables in $Z$ is an equality of the form $W=1$, where $W\in F\ast G$. If $W=z_{i_{1}}^{\varepsilon_{1}}g_{1}\cdots z_{i_{k}}^{\varepsilon_{k}}g_{k}$, with $z_{i_{j}}\in Z$, $\varepsilon_{j}=\pm 1$, and $g_{j}\in G$, then we refer to $\{z_{i_{1}},\ldots,z_{i_{k}}\}$ as the set of variables and to $\{g_{1},\ldots,g_{k}\}$ as the set of constants (or coefficients) of $W$. We occasionally write $W(z_{1},\ldots,z_{k})$ or $W(z_{1},\ldots,z_{k};g_{1},\ldots,g_{k})$ to indicate that the variables in $W$ are precisely $z_{1},\ldots,z_{k}$ and (in the latter case) the constants are precisely $g_{1},\ldots,g_{k}$. A solution for an equation $W(z_{1},\ldots,z_{k})=1$ over $G$ is an assignment for variables $z_{1},\ldots,z_{k}$ that makes $W=1$ true.
@@ 1.2. Constrained equations
Recently, constrained equations in groups have attracted significant attention, with most research focusing on free groups, hyperbolic groups, right-angled Artin groups, virtually abelian groups, and certain classes of finite groups. For a survey of recent results in this area see [7].
@@ 1.3. Spherical equations
$\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}$ $=1$ $m\geq 1,$ (1)
$\prod_{i=1}^{g}[x_{i},y_{i}]\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}$ $=1$ $g\geq 1,m\geq 0,$ (2)
$\prod_{i=1}^{g}x_{i}^{2}\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}$ $=1$ $g\geq 1,m\geq 0.$ (3)
@@ 1.5. Model of computation
2. Preliminaries: subset sum problem in groups
@@
$g=g_{1}^{\varepsilon_{1}}\cdots g_{k}^{\varepsilon_{k}}$ (5)
@@
By [30, Proposition 2.5], computational properties of ${\mathbf{SSP}}$ do not depend on the choice of a finite generating set $X$ and, hence, the problem can be abbreviated as ${\mathbf{SSP}}(G)$. The same paper, [30], provides a variety of examples of groups with ${\mathbf{NP}}$-complete (or polynomial time) subset sum problems.
Consider the infinitely generated group ${\mathbb{Z}}_{3}^{\omega}$ whose elements can be formally viewed as functions $f\colon{\mathbb{N}}\to{\mathbb{Z}}_{3}$ with finite support. For algorithmic purposes, we assume that elements of ${\mathbb{Z}}_{3}^{\omega}$ are encoded by finite ternary strings (as in [34, Section 4]). [35, Proposition 2.1] proves that ${\mathbf{SSP}}({\mathbb{Z}}_{3}^{\omega})$ is ${\mathbf{NP}}$-complete, which can be reformulated as follows.
Proposition 1 ([35, Proposition 2.1]).
@@ 3.2. Short integer solution problem
$\left\{\begin{array}[]{l}A{\overline{x}}\equiv_{p}{\overline{0}},\\ {\overline{x}}\mbox{ is short and nontrivial},\\ \end{array}\right.$ (7)
@@
Randomized ${\mathbf{SIS}}$ problem. For a matrix $A\in{\mathbb{Z}}_{p}^{n\times m}$ sampled uniformly randomly, find ${\overline{x}}\in{\mathbb{Z}}_{p}^{m}$ such that $A{\overline{x}}={\overline{0}}$, and ${\overline{x}}$ is short and nontrivial.
Theorem 2 (Goldreich, Goldwasser, Halevi, [15, Theorem 1], cf. Ajtai, [3, Theorem 1]).
@@
$n\log(p)<m<\tfrac{p}{2n^{4}}\ \mbox{ and }\ p=O(n^{c})\mbox{ for some }c>0,$ (8)
@@
Randomized ${\mathbf{ISIS}}$ problem. For a matrix $A\in{\mathbb{Z}}_{p}^{n\times m}$ and a vector ${\overline{y}}\in{\mathbb{Z}}_{p}^{n}$ sampled uniformly randomly, find ${\overline{x}}\in{\mathbb{Z}}_{p}^{m}$ such that $A{\overline{x}}={\overline{y}}$, and ${\overline{x}}$ is short.
Corollary 3.
The statement of Theorem 2 holds for the randomized ${\mathbf{ISIS}}_{\{0,1\}}$ problem.
Proof 3.1.
@@
This gives a PPT algorithm that solves ${\mathbf{SIS}}_{\{0,1\}}$ with probability at least $n^{-c_{0}}$.
4. Spherical equations over $G_{p,n}$
@@
Proof 4.1.
Straightforward verification.
4.1. Spherical equations over $G_{p,n}$: generic-case
Here, we investigate generic-case hardness for spherical equations, i.e., hardness of a “typical” equation, see [31, Chapter 10] for basic definitions of generic-case complexity.
Proposition 5.
@@ 4.1. Spherical equations over $G_{p,n}$: generic-case
is a solution that can be computed in polynomial time.
@@
that converges to 00 exponentially fast. The obtained bound $\frac{1}{2^{s/2}}$ does not depend on $p$ or $n$. Therefore, the bound works on the whole ${\mathcal{I}}_{s}$.
Lemma 7.
@@
Proof 4.4.
By Lemma 6, we may assume that $m\geq s/2$, in which case
@@
which converges to $1$ exponentially fast as $s\to\infty$.
Corollary 8.
@@
Proof 4.5.
Follows from Proposition 5 and Lemma 7
@@
$C_{p,n}=\left\{\,({\overline{c}},1)\;\middle|\;{\overline{c}}\in{\mathbb{Z}}_{p}^{n}\,\right\}.$ (9)
From now on we assume that all $c_{i}\in C_{p,n}$. In that case condition (S1) of Lemma 4 is trivially satisfied and condition (S2) translates into the following:
$\exists\alpha_{i}\in{\mathbb{Z}}_{p}^{\ast}\ \ \mbox{ s.t. }\ \sum_{i=1}^{m}\alpha_{i}^{-1}{\overline{y}}_{i}\equiv_{p}{\overline{0}}.$ (10)
@@
4.2. Spherical equations over $G_{p,n}$: worst case
@@
$\prod_{i}^{m}z_{i}^{-1}c_{{\overline{v}}_{i}}z_{i}=c_{{\overline{v}}^{\prime}}.$ (11)
Proposition 9.
@@
Proof 4.7.
By Proposition 9, $I\to E_{I}$ is a many-one polynomial-time reduction of ${\mathbf{SSP}}({\mathbb{Z}}_{3}^{n})$ to spherical equations over ${\mathbf{SSP}}({\mathbb{Z}}_{3,n}\rtimes{\mathbb{Z}}_{3}^{\ast})$. By Proposition 1, ${\mathbf{SSP}}$ is ${\mathbf{NP}}$-complete for groups $\{{\mathbb{Z}}_{3}^{n}\}_{n=1}^{\infty}$. Hence the result.
@@ 5. Constrained spherical equations: average-case hardness
In this section we discuss spherical equations over the class of groups ${\mathcal{G}}=\{G_{p,n}\}$ with constraints on values of $z_{i}$’s. In full generality the problem can be formulated as follows. Given a group $G\in{\mathcal{G}}$, a spherical equation $\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=1$ over $G$, and subsets $Z_{1},\dots,Z_{m}\subseteq G$, find $z_{1},\dots,z_{m}\in G$ satisfying
@@
It is not hard to prove that ${\mathbf{CISE}}$ for groups $G_{p,n}$ is hard in the worst case. In fact, for $p=3$, the proof of Proposition 9 works as is, with constraints $Z_{i}=\{({\overline{0}},1),({\overline{0}},2)\}$ (so chosen sets $Z_{i}$ effectively define no constraints). Using these constraints we can easily achieve the same result for any odd value of $p$.
Theorem 11.
@@
Proof 5.1.
For a given instance ${\overline{v}}_{1},\dots,{\overline{v}}_{m},{\overline{v}}\in{\mathbb{Z}}_{p}^{n}$, call it $I$, of ${\mathbf{SSP}}({\mathbb{Z}}_{p}^{n})$, as in the proof of Proposition 9, construct a constrained inhomogeneous spherical equation $E_{I}$ over ${\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$
@@
and observe that $I$ is a positive instance of ${\mathbf{SSP}}({\mathbb{Z}}_{p}^{n})$ if and only if $E_{I}$ has a solution. This gives a polynomial time reduction from ${\mathbf{SSP}}$ over groups $\{{\mathbb{Z}}_{p}^{n}\}_{n=1}^{\infty}$ to decidability of constrained inhomogeneous spherical equations over ${\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$.
@@
$\left\{\begin{array}[]{l}\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=c\ \ \mbox{ with }c_{i},c\in C_{p,n}\\ z_{i}=({\overline{0}},1)\mbox{ or }({\overline{0}},2^{-1}).\end{array}\right.$ (12)
@@
$\left\{\begin{array}[]{l}\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=c\ \ \mbox{ with }c_{i},c\in C_{p,n}\\ z_{i}=({\overline{0}},1),({\overline{0}},2^{-1}),\mbox{ or }({\overline{0}},3^{-1}).\end{array}\right.$ (13)
@@
This establishes a one-to-one correspondence between ${\mathbf{ISIS}}_{\{0,1\}}$ and ${\mathbf{CISE}}_{\{1,2\}}$ that translates the uniform distribution for the instances of ${\mathbf{ISIS}}_{\{0,1\}}$ to the uniform distribution for the instances of ${\mathbf{CISE}}_{\{1,2\}}$ (the element ${\overline{v}}^{\prime}\in C_{p,n}$ is uniformly distributed because ${\overline{y}}$ uniformly distributed). A similar reduction can be used to establish a one-to-one correspondence between ${\mathbf{SIS}}_{\{-1,0,1\}}$ and ${\mathbf{CISE}}_{\{1,2,3\}}$.
Remark 5.3.
The one-to-one correspondence established in the proof of Theorem 12 also implies that uniformly chosen instances of ${\mathbf{CISE}}_{\{1,2\}}$ and ${\mathbf{CISE}}_{\{1,2,3\}}$ have solutions with probability approaching $1$ as $n\to\infty$.
Remark 5.4.